OpenStudy (anonymous):

Use the ZPP (Zero Product Property) to solve the equation. Check your answer. Show all your work. x(x+3)=18

OpenStudy (tkhunny):

In order to use the ZPP, you should rearrange so that you have zero, no?

OpenStudy (anonymous):

Right, so then the new equation would become x(x+3)-18=0?

OpenStudy (tkhunny):

Good. No, expand and refactor in the new condition.

OpenStudy (anonymous):

Can you help me with the next step... because I'm trying to plug in numbers into the equation and so it becomes 18-18=0. Sorry.

OpenStudy (tkhunny):

Why are you in "plug in" mode? Did you learn factoring? Expand, first! You need to factor or use the quadratic formula.

OpenStudy (tkhunny):

x(x+3)-18=0 \(x^{2} + 3x - 18 = 0\) Looking familiar?

OpenStudy (anonymous):

Yeah, but my teacher wants me to ZPP it. So like for example: (x+1)(x+3)=0 1. (-1+1)(-1+3)=0 0 (2) = 0 0 = 0 2. (-3+1)(-3+3)=0 (-2)0=0 0=0 And so the solutions would be -1 and -3. I don't understand why she wants us to do it this way....but those are the instructions. I understand how to factor and to use the quadratic formula....however I don't know if this is how she would like for me to do the problem.

OpenStudy (tkhunny):

If you know how to factor, factor this: \(x^{2} + 3x - 18\)

OpenStudy (anonymous):

Yeah, but my teacher wants me to ZPP it. So like for example: (x+1)(x+3)=0 1. (-1+1)(-1+3)=0 0 (2) = 0 0 = 0 2. (-3+1)(-3+3)=0 (-2)0=0 0=0 And so the solutions would be -1 and -3. I don't understand why she wants us to do it this way....but those are the instructions. I understand how to factor and to use the quadratic formula....however I don't know if this is how she would like for me to do the problem.

OpenStudy (tkhunny):

So, you're not going to factor it? Why do you keep telling my the same thing? FACTOR IT!!!!!